The length and width of a rectangle are in a 4:7 ratio. The perimeter of the rectangle is 352. What are the length and width?
Answer by TimothyLamb(4379)

L/w = 4/7
L = (4/7)w
---
2L + 2w = 352
2(4/7)w + 2w = 352
(8/7)w + (14/7)w = 352
(22/7)w = 352
w = 352*(7/22)
w = 112
L = (4/7)w
L = (4/7)*112
L = 64
---
answer:
w = 112
L = 64
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php